OpenStudy (anonymous):

Which mathematical sentence represents the solution for a in the equation 2b = bc + 5a?

OpenStudy (anonymous):

is there any answer choices?

OpenStudy (anonymous):

options a=2b-bc a=2b-bc/5 a=2b-c/5 a=2b+bc/5

OpenStudy (anonymous):

a equals two b minus b c all over five answer is c

OpenStudy (anonymous):

thank you :D i was so confused

OpenStudy (anonymous):

your welcome no problem